OpenStudy (anonymous):

what is the geometric mean of 3 and 7?

OpenStudy (anonymous):

To find the mean of something, you add the terms and then divide by the amount of them. So, they equal 10 and there are two of them. 10/2 = 5. The mean should be 5.

OpenStudy (anonymous):

Square root of product

OpenStudy (anonymous):

I thought I multiply the two and find the square root?

OpenStudy (anonymous):

I repeat Square root of product

OpenStudy (anonymous):

Don't listen to me... O_O I may very well be wrong.

OpenStudy (anonymous):

so I should square root 21?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

but that just gets a weird number

OpenStudy (anonymous):

If it was 3 numbers, then cube root of 3 multiplied together....

OpenStudy (anonymous):

ah i see

OpenStudy (anonymous):

It's just a different kind of average....

OpenStudy (anonymous):

So would I do something like 2*square root of the product?

OpenStudy (anonymous):

Nope, just sqrt 21, that's the GM

OpenStudy (anonymous):

so I just leave it at square root 21?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

why's that?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

why do i leave it at \[\sqrt{21}\] ?

OpenStudy (anonymous):

Is there something wrong with sqrt 21 ?

OpenStudy (anonymous):

don't I have to actually sqaure root it and make it X *\[\sqrt{X}\]

OpenStudy (anonymous):

or something like that?

OpenStudy (anonymous):

I don't understand, the square root of 21 is roughly 4.5825, unless you are asked to give a decimal approximation, just leave it as sqrt 21

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

ur welcome